OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0.

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

No .

OpenStudy (anonymous):

Wait, With The Box & Diamond Problem ?

OpenStudy (lgbasallote):

i dont think that's factorable

OpenStudy (anonymous):

How Do You Solve It Then ?

OpenStudy (anonymous):

quadratic formula. -.- it's so much more complicated than factoring.

OpenStudy (lgbasallote):

\[\LARGE x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] are you familiar with this? @Briannaw

OpenStudy (anonymous):

Yes I Am Can You Help Me Set It Up ?

OpenStudy (anonymous):

quadratic equation: \(\large ax^2+bx+c=0 \) your equation: \(\large x^2+9x+9=0 \) first you need to identify these values a= b= c= can you do that?

OpenStudy (anonymous):

a= 1 b=9 c=9

OpenStudy (anonymous):

ok.. now all you have to do is compute: \[\large x=\frac{-(9)\pm\sqrt{(9)^2-4(1)(9)}}{2(1)} \]